OpenStudy (anonymous):

Claire has received scores of 85, 88, 87, and 80 on her algebra tests. What score must she receive on the fifth test to have an overall test score average of at least 82? A) 70 or greater B) 68 or greater C) 71 or greater D) 69 or greater

OpenStudy (anonymous):

(85+88+87+80 + X) / 5 >= 82 (340+X) / 5 >= 82 340+X >= 410 x >= 70

OpenStudy (anonymous):

thank you so much!